OpenStudy (anonymous):

let ∫(3,0) f(x)dx=6. what is the average value of f(x) on the interval x=0 to x=3?

OpenStudy (anonymous):

adamaero can you help?

OpenStudy (adamaero):

You should tag someone else (with @). I'm not sure what it's asking. http://www.math.msu.edu/~robertbe/riemann.pdf

OpenStudy (anonymous):

∫ the 3 is on top and the 0 is on bottom of this sybol, i just dont know how to type it like that

OpenStudy (adamaero):

\[\int\limits_{0}^{3}\]

OpenStudy (anonymous):

\[\int\limits_{0}^{3} f(x)dx=6. what is the average value of f(x) on the interval x=0 \to x=3?\]

OpenStudy (anonymous):

\[\int\limits_{0}^{3} f(x)dx=6\]

OpenStudy (anonymous):

Given that, i need to find the average value of f(x) over the interval x=0 to x=3